# PEAN: A Diffusion-Based Prior-Enhanced Attention Network for Scene Text Image Super-Resolution

### v2

@@ Abstract.
Scene text image super-resolution (STISR) aims at simultaneously increasing the resolution and readability of low-resolution scene text images, thus boosting the performance of the downstream recognition task. Two factors in scene text images, visual structure and semantic information, affect the recognition performance significantly. To mitigate the effects from these factors, this paper proposes a Prior-Enhanced Attention Network (PEAN). Specifically, an attention-based modulation module is leveraged to understand scene text images by neatly perceiving the local and global dependence of images, despite the shape of the text. Meanwhile, a diffusion-based module is developed to enhance the text prior, hence offering better guidance for the SR network to generate SR images with higher semantic accuracy. Additionally, a multi-task learning paradigm is employed to optimize the network, enabling the model to generate legible SR images. As a result, PEAN establishes new SOTA results on the TextZoom benchmark. Experiments are also conducted to analyze the importance of the enhanced text prior as a means of improving the performance of the SR network. Code will be made available at https://github.com/jdfxzzy/PEAN.
@@
ccs: Computing methodologies Reconstruction
@@ 1. Introduction
Despite improved performance achieved by these approaches, the dominance of visual structure and semantic information persists, as two critical issues in previous studies remain unresolved. Firstly, previous STISR methods (Chen et al., 2021, 2022; Ma et al., 2022; Wang et al., 2020; Zhao et al., 2022) rely on Sequential Residual Blocks (SRB) to extract visual features. This module, containing several CNN-BiLSTM layers, has difficulty in restoring the complete visual structure of images containing long or deformed text string due to its inherent demerits, i.e., the performance bottleneck of capturing long-range dependence (Dosovitskiy et al., 2020; Raghu et al., 2021). Secondly, the introduction of the primary text prior, originating from the interference of low-quality images on recognizers, prevents the SR network from generating images that contain correct semantic information. Recently, C3-STISR (Zhao et al., 2022) has employed a language model (Fang et al., 2021a) into STISR, utilizing its learned linguistic knowledge to rectify the text prior. Although the rectified prior demonstrates some effectiveness, it lacks sufficient strength in guiding the SR network to produce images with high semantic accuracy.
@@
We propose a Prior-Enhanced Attention Network (PEAN) to tackle issues caused by the two factors. To begin with, an Attention-based Modulation Module (AMM) is proposed to substitute the SRB, endowing the network with a larger receptive field to images, thereby restoring the visual structure of images with text in various shapes and lengths (shown in Figure 1(d)). Horizontal and vertical strip-wise attention mechanisms (Dong et al., 2022; Huang et al., 2023; Tsai et al., 2022) are employed in AMM. Among them, the horizontal attention mechanism can capture the dependence between characters while the vertical attention mechanism can capture the structural information within a character (Zhao et al., 2021). However, the lack of semantic information limits the capability of such model. As demonstrated by previous works (Dong et al., 2016; Dahl et al., 2017), leveraging strong prior information to restrict the solution space plays a vital role in SR problems. Notably, the text prior derived from high-resolution (HR) images is a robust choice for STISR, in view of the high recognition accuracy of HR images. Consequently, we conduct an exploratory experiment wherein we substitute the text prior from LR images (TP-LR) with the text prior from HR images (TP-HR) within such model, yielding superior outcomes (see Figures 1(e) and (f) for comparison, details can be found in § 3). This inspires the design of a module for enhancing the primary text prior, resulting in the creation of the Enhanced Text Prior (ETP), which is comparable in effectiveness to TP-HR (shown in Figures 1(f) and (g)). The ETP provides valuable guidance to the SR network, promoting the generation of SR images with high semantic accuracy. Given the remarkable performance of diffusion models (Ho et al., 2020; Song et al., 2021), we propose a diffusion-based Text Prior Enhancement Module (TPEM) to obtain the ETP owing to their ability to map complex distributions (Xia et al., 2023). In addition, considering that the goal of STISR is to increase the resolution and readability of LR scene text images, we adopt the Multi-Task Learning (MTL) paradigm in the training phase, where the image restoration task aims at generating high-quality SR images, and the text recognition task stimulates the model to generate more readable SR results. In a nutshell, main contributions of our work are three-fold:
@@ 2. Related Work
In this section, we give a brief review of recent progress in scene text image super-resolution, scene text recognition and diffusion models, which are relevant to our work.
@@ 3.2. Text Prior Enhancement Module
As demonstrated by previous works (Dahl et al., 2017; Dong et al., 2016), strong prior information plays a pivotal role in solving SR problems, while the primary text prior applied in previous works is not powerful enough because it originates from LR scene text images. Our exploratory experiments, detailed in § 3, also underline the influential role of TP-HR in guiding the SR network to generate images with improved semantic accuracy for the model. Therefore, we introduce the TPEM to obtain the ETP, which can effectively guide the SR network, similarly to the efficacy of TP-HR. The core component of TPEM is the denoiser, denoted as $f_{\theta}$, which leverages the reverse diffusion process (Ho et al., 2020) to estimate the enhanced prior, providing substantial semantic guidance to the SR network.
@@ 3.2.2. Reverse Diffusion Process
In the reverse diffusion process, Gaussian noise gradually transforms into the ETP, denoted as $P^{\mathrm{e}}\in\mathbb{R}^{L\times|\mathcal{A}|}$, conditioned on the primary text prior $P^{\mathrm{l}}\in\mathbb{R}^{L\times|\mathcal{A}|}$. The latter is the output recognition probability sequence of the TPG, with $I^{\mathrm{lr}}$ as input. This process can be mathematically formulated as follows:
@@
where $f_{\theta}$ is an MLP-based denoising network, the architecture of which can be found in the Supplementary Material. Similar to previous works (Liu et al., 2023; Xia et al., 2023; Yang et al., 2023; Zhao et al., 2023), we opt to directly estimate $P^{\mathrm{e}}$ instead of $\epsilon$ for performance improvement. Experiments in § 5 verifies the effectiveness of this design. The whole process is supervised by the MAE and CTC loss (Graves et al., 2006), given by:
@@ 4.1. Datasets and Evaluation Metrics
We conduct experiments on the TextZoom (Wang et al., 2020) benchmark. However, due to some inherent drawbacks of it, these experiments are not enough to reflect that PEAN surely has the ability to restore the complete visual structure. Therefore, we select 651 images with the resolution no greater than $16\times 64$ as LR images from the IIIT5K (Mishra et al., 2012), SVTP (Phan et al., 2013) and IC15 (Karatzas et al., 2015) datasets for evaluation. Details of the datasets and drawbacks of TextZoom can be found in the Supplementary Material. For TextZoom, following previous works (Wang et al., 2020; Chen et al., 2021, 2022; Ma et al., 2023, 2022; Zhao et al., 2022; Guo et al., 2023; Zhang et al., 2023), we adopt ASTER (Shi et al., 2018), MORAN (Luo et al., 2019) and CRNN (Shi et al., 2016a) for evaluation. In the Supplementary Material, we also adopt three recent Transformer-based recognizers, namely, MGP-STR (Wang et al., 2022), ABINet (Fang et al., 2021a) and VisionLAN (Wang et al., 2021) for evaluation. For the constructed dataset, PSNR and SSIM (Wang et al., 2004) are selected as metrics to evaluate the image quality.
@@ 4.2. Implementation Details
Our model is implemented with the Pytorch 1.10 deep learning library (Paszke et al., 2019). All of the experiments are conducted on 1 NVIDIA TITAN RTX GPU. In the training phase, the model is trained for 200 epochs and optimized by the AdamW (Loshchilov and Hutter, 2019) optimizer. The learning rate and the size of the mini-batch are set as 0.001 and 32 respectively. The weight of each loss is set as $\lambda_{1}=1$, $\lambda_{2}=1$, $\lambda_{3}=0.8$, $\lambda_{4}=75$, $\lambda_{5}=1$. In terms of the network architecture, the number of blocks in AMM is 6. The sampling timestep in the TPEM is set as 1. Following CRNN (Shi et al., 2016a), the number of convolutional layers applied in the ARM is 6. We first drop the TPEM and pre-train the model with TP-HR, then the TPEM is introduced and weights of parameters obtained by pre-training are initialized to continue the fine-tuning process of the model. Of note, if this setting is abandoned, PEAN can still achieve SOTA performance on the TextZoom (Wang et al., 2020) dataset. Ablation studies about the weight of loss functions and the pre-training and fine-tuning setting can be found in the Supplementary Material. In the main paper, we study our method on top of PARSeq (Bautista and Atienza, 2022) as the TPG. In the Supplementary Material, we also study the cases when CRNN (Shi et al., 2016a) and ABINet (Fang et al., 2021a) are adopted as the TPG.
@@ 4.4.1. Text Prior and the Enhancement Module
Since previous works such as Zhao et al. (Zhao et al., 2022) introduce a Language Model (LM) (Fang et al., 2021a) to rectify the text prior, here we compare the TPEM with several LMs used in other tasks of scene text images. Among them, GSRM is an autoregressive-based LM proposed by SRN (Yu et al., 2020), while PD is the parallel Transformer decoder in PIMNet (Qiao et al., 2021) based on an easy-first decoding strategy. BCN (Fang et al., 2021a) is the autoencoding-based LM applied in C3-STISR (Zhao et al., 2022), proposed in ABINet (Fang et al., 2021a). Experimental results are shown in Table 4, from which we can conclude that the employment of the diffusion-based TPEM brings more performance gain compared with other variants. This can be attributed to the powerful distribution mapping capability (Xia et al., 2023) of diffusion models.
Table 4. Ablation study on different enhancement modules.
  Methods Easy Medium Hard Average
GSRM (Yu et al., 2020) 76.3 58.8 40.7 59.7
PD (Qiao et al., 2021) 78.1 61.6 43.4 62.1
BCN (Fang et al., 2021a) 79.6 61.9 44.8 63.2
PEAN 84.5 71.4 52.9 70.6
4.4.2. Optimization Objective for Training the TPEM
Regarding to most of the diffusion-based models (Saharia et al., 2023; Gao et al., 2023a; Ren et al., 2023; Rombach et al., 2022), researchers typically constrain the estimated noise (denoted as $\epsilon$) through the loss function, as supported by Ho et al. (Ho et al., 2020), for the sake of improving sample quality. However, as discussed in § 3.2.2, in the context of the TPEM, we choose to directly estimate the ETP (denoted as $P^{\mathrm{e}}$) instead of focusing on the noise for enhanced performance. In this regard, we conduct experiments with both optimization objectives, and the results are presented in Table 5. We argue that selecting $P^{\mathrm{e}}$ as the optimization target can yield superior performance in our proposed PEAN because it encourages the network to concentrate on the enhancement of the text prior itself. This approach also opens the possibility of introducing additional constraints on the ETP, such as the CTC loss (Graves et al., 2006) denoted as $\mathcal{L}_{\mathrm{ctc}}^{\mathrm{t}}$.
Table 5. Ablation study about the objective of optimization.
  Objectives Easy Medium Hard Average
$\epsilon$ 80.0 65.6 46.3 65.0
$P^{\mathrm{e}}$ 84.5 71.4 52.9 70.6
@@ 4.4.3. Impact of the AMM
Here we perform experiments to justify the superiority of the AMM adopted in our model. As presented in Table 6, the comparison reveals the following points: (1) Our proposed AMM handles the STISR task in a more effective way than the CNN-BiLSTM-based SRB (Wang et al., 2020). The local inductive bias of the CNN (Dosovitskiy et al., 2020; Raghu et al., 2021) and the rigid nature of BiLSTM (Vaswani et al., 2017; Zhu et al., 2023b) limit its performance in recovering the visual structure of images with long or deformed text (Chen et al., 2021; Ma et al., 2022), while the AMM can solve this problem. (2) The typical ViT-based architectures (Dong et al., 2022; Dosovitskiy et al., 2020; Liu et al., 2021) show trivial performance, as a result of the tendentious design for high-level vision tasks, which neglects the inherent characteristic of STISR. (3) Although employing the strip-wise attention mechanism, Stripformer (Tsai et al., 2022), which also leverages the conditional positional encodings (Chu et al., 2023) and several residual blocks for the image deblurring task, fails in the STISR task. On average, PEAN obtains an improvement of the performance of +5.9 from the AMM. Further experiments in the Supplementary Material show that even without the MTL paradigm, the AMM still outperforms the SRB (Wang et al., 2020).
@@ 4.4.4. Effect of the MTL Paradigm
In this part, we conduct experiments to show the effect of the MTL paradigm introduced in our work. The results shown in Table 7 indicate that: (1) The stroke-based SFM loss (Chen et al., 2022) provides an average contribution of +4.6 for the improved performance, which convinces us that the preservation of visual structure plays a vital role in STISR. (2) The text recognition task employed in our work is also indispensable. The CTC (Graves et al., 2006) loss applied on the output of ARM ($\,\mathcal{L}_{\mathrm{ctc}}^{\mathrm{a}}$) brings correct semantic constraint for the AMM. Besides, the CTC loss imposed on the ETP ($\,\mathcal{L}_{\mathrm{ctc}}^{\mathrm{t}}$) ensures the coherence between the text prior and the ground truth text label, guiding the training of the enhancement module in a precise way. We provide more detailed experiments in the Supplementary Material.

### v3

@@ Abstract.
Scene text image super-resolution (STISR) aims at simultaneously increasing the resolution and readability of low-resolution scene text images, thus boosting the performance of the downstream recognition task. Two factors in scene text images, visual structure and semantic information, affect the recognition performance significantly. To mitigate the effects from these factors, this paper proposes a Prior-Enhanced Attention Network (PEAN). Specifically, an attention-based modulation module is leveraged to understand scene text images by neatly perceiving the local and global dependence of images, despite the shape of the text. Meanwhile, a diffusion-based module is developed to enhance the text prior, hence offering better guidance for the SR network to generate SR images with higher semantic accuracy. Additionally, a multi-task learning paradigm is employed to optimize the network, enabling the model to generate legible SR images. As a result, PEAN establishes new SOTA results on the TextZoom benchmark. Experiments are also conducted to analyze the importance of the enhanced text prior as a means of improving the performance of the SR network. Code is available at https://github.com/jdfxzzy/PEAN.
@@
copyright: acmlicensedjournalyear: 2024doi: 10.1145/3664647.3680974conference: Proceedings of the 32nd ACM International Conference on Multimedia; October 28–November 1, 2024; Melbourne, VIC, Australia.booktitle: Proceedings of the 32nd ACM International Conference on Multimedia (MM ’24), October 28–November 1, 2024, Melbourne, VIC, Australiaisbn: 979-8-4007-0686-8/24/10ccs: Computing methodologies Reconstruction
@@ 1. Introduction
Despite improved performance achieved by these approaches, the dominance of visual structure and semantic information persists, as two critical issues in previous studies remain unresolved. Firstly, previous STISR methods (Chen et al., 2021, 2022; Ma et al., 2022; Wang et al., 2020; Zhao et al., 2022, 2021) rely on Sequential Residual Blocks (SRB) to extract visual features. This module, containing several CNN-BiLSTM layers, has difficulty in restoring the complete visual structure of images containing long or deformed text string due to its inherent demerits, i.e., the performance bottleneck of capturing long-range dependence (Dosovitskiy et al., 2020; Raghu et al., 2021). Secondly, the introduction of the primary text prior, originating from the interference of low-quality images on recognizers, prevents the SR network from generating images that contain correct semantic information. Recently, C3-STISR (Zhao et al., 2022) has employed a language model (Fang et al., 2021a) into STISR, utilizing its learned linguistic knowledge to rectify the text prior. Although the rectified prior demonstrates some effectiveness, it lacks sufficient strength in guiding the SR network to produce images with high semantic accuracy.
@@
We propose a Prior-Enhanced Attention Network (PEAN) to tackle issues caused by the two factors. To begin with, an Attention-based Modulation Module (AMM) is proposed to substitute the SRB, endowing the network with a larger receptive field to images, thereby restoring the visual structure of images with text in various shapes and lengths (shown in Figure 1(d)). Horizontal and vertical strip-wise attention mechanisms (Dong et al., 2022; Huang et al., 2023; Tsai et al., 2022) are employed in AMM. Among them, the horizontal attention mechanism can capture the dependence between characters while the vertical attention mechanism can capture the structural information within a character (Zhao et al., 2021). However, the lack of semantic information limits the capability of such model. As demonstrated by previous works (Dong et al., 2016; Dahl et al., 2017), leveraging strong prior information to restrict the solution space plays a vital role in SR problems. Notably, the text prior derived from high-resolution (HR) images is a robust choice for STISR, in view of the high recognition accuracy of HR images. Consequently, we conduct an exploratory experiment wherein we substitute the text prior from LR images (TP-LR) with the text prior from HR images (TP-HR) within such model, yielding superior outcomes (see Figures 1(e) and (f) for comparison, details can be found in § 4.4.1). This inspires the design of a module for enhancing the primary text prior, resulting in the creation of the Enhanced Text Prior (ETP), which is comparable in effectiveness to TP-HR (shown in Figures 1(f) and (g)). The ETP provides valuable guidance to the SR network, promoting the generation of SR images with high semantic accuracy. Given the remarkable performance of diffusion models (Ho et al., 2020; Song et al., 2021), we propose a diffusion-based Text Prior Enhancement Module (TPEM) to obtain the ETP owing to their ability to map complex distributions (Xia et al., 2023). In addition, considering that the goal of STISR is to increase the resolution and readability of LR scene text images, we adopt the Multi-Task Learning (MTL) paradigm in the training phase, where the image restoration task aims at generating high-quality SR images, and the text recognition task stimulates the model to generate more readable SR results. In a nutshell, main contributions of our work are three-fold:
@@ 3.2. Text Prior Enhancement Module
As demonstrated by previous works (Dahl et al., 2017; Dong et al., 2016), strong prior information plays a pivotal role in solving SR problems, while the primary text prior applied in previous works is not powerful enough because it originates from LR scene text images. Our exploratory experiments, detailed in § 4.4.1, also underline the influential role of TP-HR in guiding the SR network to generate images with improved semantic accuracy for the model. Therefore, we introduce the TPEM to obtain the ETP, which can effectively guide the SR network, similarly to the efficacy of TP-HR. The core component of TPEM is the denoiser, denoted as $f_{\theta}$, which leverages the reverse diffusion process (Ho et al., 2020) to estimate the enhanced prior, providing substantial semantic guidance to the SR network.
@@ 3.2.2. Reverse Diffusion Process
In the reverse diffusion process, Gaussian noise gradually transforms into the ETP, denoted as $P^{\mathrm{e}}\in\mathbb{R}^{L\times|\mathcal{A}|}$, conditioned on the primary text prior $P^{\mathrm{l}}\in\mathbb{R}^{L\times|\mathcal{A}|}$. The latter is the output recognition probability sequence of the TPG, with $I^{\mathrm{lr}}$ as input. This process can be formulated as follows:
@@
where $f_{\theta}$ is an MLP-based denoising network, the architecture of which can be found in § A of the Supplementary Material. Similar to previous works (Liu et al., 2023; Xia et al., 2023; Yang et al., 2023; Zhao et al., 2023), we opt to directly estimate $P^{\mathrm{e}}$ instead of $\epsilon$ for performance improvement. Experiments in § 3 verifies the effectiveness of this design. The whole process is supervised by the MAE and CTC loss (Graves et al., 2006), given by:
@@ 4.1. Datasets and Evaluation Metrics
We conduct experiments on the TextZoom (Wang et al., 2020) benchmark. However, due to some inherent drawbacks of it, these experiments are not enough to reflect that PEAN surely has the ability to restore the complete visual structure. Therefore, we select 651 images with the resolution no greater than $16\times 64$ as LR images from the IIIT5K (Mishra et al., 2012), SVTP (Phan et al., 2013) and IC15 (Karatzas et al., 2015) datasets for evaluation. Details of the datasets and drawbacks of TextZoom can be found in § B.2 of the Supplementary Material. For TextZoom, following previous works (Wang et al., 2020; Chen et al., 2021, 2022; Ma et al., 2023, 2022; Zhao et al., 2022; Guo et al., 2023; Zhang et al., 2023), we adopt ASTER (Shi et al., 2018), MORAN (Luo et al., 2019) and CRNN (Shi et al., 2016a) for evaluation. In § B.1 of the Supplementary Material, we also adopt three recent Transformer-based recognizers, i.e., MGP-STR (Wang et al., 2022), ABINet (Fang et al., 2021a) and VisionLAN (Wang et al., 2021) for evaluation. For the constructed dataset, PSNR and SSIM (Wang et al., 2004) are selected as metrics to evaluate the image quality.
@@ 4.2. Implementation Details
Our model is implemented with the Pytorch 1.10 deep learning library (Paszke et al., 2019). All of the experiments are conducted on 1 NVIDIA TITAN RTX GPU. In the training phase, the model is trained for 200 epochs and optimized by the AdamW (Loshchilov and Hutter, 2019) optimizer. The learning rate and the size of the mini-batch are set as 0.001 and 32 respectively. The weight of each loss is set as $\lambda_{1}=1$, $\lambda_{2}=1$, $\lambda_{3}=0.8$, $\lambda_{4}=75$, $\lambda_{5}=1$. In terms of the network architecture, the number of blocks in AMM is 6. The sampling timestep in the TPEM is set as 1. Following CRNN (Shi et al., 2016a), the number of convolutional layers applied in the ARM is 6. We first drop the TPEM and pre-train the model with TP-HR, then the TPEM is introduced and weights of parameters obtained by pre-training are initialized to continue the fine-tuning process of the model. Of note, if this setting is abandoned, PEAN can still achieve SOTA performance on the TextZoom (Wang et al., 2020) dataset. Ablation studies about the weight of loss functions and the pre-training and fine-tuning setting can be found in § XV and § XVIII of the Supplementary Material respectively. In the main paper, we study our method on top of PARSeq (Bautista and Atienza, 2022) as the TPG. In § XVIII of the Supplementary Material, we also study the cases when CRNN (Shi et al., 2016a) and ABINet (Fang et al., 2021a) are adopted as the TPG.
@@ 4.4.2. Impact of the AMM
Here we perform experiments to justify the superiority of the AMM adopted in our model. As presented in Table 4, the comparison reveals the following points: (1) Our proposed AMM handles the STISR task in a more effective way than the CNN-BiLSTM-based SRB (Wang et al., 2020). The local inductive bias of the CNN (Dosovitskiy et al., 2020; Raghu et al., 2021) and the rigid nature of BiLSTM (Vaswani et al., 2017; Zhu et al., 2023b) limit its performance in recovering the visual structure of images with long or deformed text (Chen et al., 2021; Ma et al., 2022), while the AMM can solve this problem. (2) The typical ViT-based architectures (Dong et al., 2022; Dosovitskiy et al., 2020; Liu et al., 2021) show trivial performance, as a result of the tendentious design for high-level vision tasks, which neglects the inherent characteristic of STISR. (3) Although employing the strip-wise attention mechanism, Stripformer (Tsai et al., 2022), which also leverages the conditional positional encodings (Chu et al., 2023) and several residual blocks for the image deblurring task, fails in the STISR task. On average, PEAN obtains an improvement of the performance of +5.9 from the AMM. Experiments in § XIII of the Supplementary Material show that even without the MTL paradigm, the AMM still outperforms the SRB (Wang et al., 2020).
@@ 4.4.3. Effect of the MTL Paradigm
In this part, we conduct experiments to show the effect of the MTL paradigm introduced in our work. The results shown in Table 5 indicate that: (1) The stroke-based SFM loss (Chen et al., 2022) provides an average contribution of +4.6 for the improved performance, which convinces us that the preservation of visual structure plays a vital role in STISR. (2) The text recognition task employed in our work is also indispensable. The CTC (Graves et al., 2006) loss applied on the output of ARM ($\,\mathcal{L}_{\mathrm{ctc}}^{\mathrm{a}}$) brings correct semantic constraint for the AMM. Besides, the CTC loss imposed on the ETP ($\,\mathcal{L}_{\mathrm{ctc}}^{\mathrm{t}}$) ensures the coherence between the text prior and the ground truth text label, guiding the training of the enhancement module in a precise way. We provide more detailed experiments in § XIII of the Supplementary Material.